OpenStudy (anonymous):

Consider matrix A (1 2 1) (6 -1 0) (-1 -2 -1) Find the eigenvalues and the eigenvectors. I already found the eigenvalues to be -4, 0, 3. I am having trouble however finding the eigenvectors. Can someone show me how to solve for the eigenvectors? Any help would be appreciated.

OpenStudy (anonymous):

To get an eigenvector, you solve for \(\overrightarrow{x}\) in \((\lambda I-A)\overrightarrow{x}=\overrightarrow{0}\). \(I\) is the identity matrix and \(A\) is your matrix

OpenStudy (anonymous):

So to find the corresponding eigenvector for -4, we have: \[ \left( \begin{bmatrix} -4 & 0 & 0 \\ 0 & -4 & 0 \\ 0 & 0 & -4 \end{bmatrix} - \begin{bmatrix} 1 & 2 & 1 \\ 6 & -1 & 0 \\ -1 & -2 & -1 \end{bmatrix} \right) \overrightarrow{x} = \overrightarrow{0} \]

OpenStudy (anonymous):

Just do matrix subtraction, then do elimination or whatever to solve for \(\overrightarrow{x}\).

OpenStudy (anonymous):

Note that eigenvectors are magnitude independent